1. Suppose we know that the total budget is $48, price of beef (PGI) İS $3.00/lb, and price of pork (Pa) is S2.00lb. (1) Graphically and mathematically express the budget line. (Make sure you find the two end points of BL. in the graph.) (2) What is the slope of your above budget line (Please use the price ratios to calculate it SHOW YOUR WORK)? (3) If PG increases to S4.00-1b, while the TB ind P a dont change, graphically and mathematically express the new BL slope of new BL? what is the (4)IFTB-s60, while price of beef (Po) and price of pork (P ac)remain $3.00l express the BL. What is the slope of BL? and $2.00/lb each, gruphically and mathematically
Answer #1

1. Total budget = $ 48,

Price of Beef, PG1 = $ 3 /lb

Price of pork, PG2 = $ 2/lb

Let x unit of Beef is purchased and y unit of pork

3x + 2y = 48

Let us assume when y= 0

No. Of units of Beef purchased = 16

Similarly pork = 24

Slope is Price of x /price of y. Here we have beef on y axis and pork on x axis.

Slope = Px/Py = 2/3

2. The slope of the budget line = Price of pork/price of Beef = 2/3

3. New budget line, 4x + 2y = 48

The two coordinates are

Beef = 12

Pork = 24

Slope = 2/4 = 1/2

4. When budget is $ 60

New budget line

3x + 2y = 60

The two coordinates are

Beef = 20

Pork = 30

Slope of budget line = 2/3

Please refer the attached picture below all budget line are drawn on same page.

047 14 /2 24 241 24 夕4 $4,

Please contact if having any query thank you.